In the space of X=E^1, what is the complement of the following:
(1,2]
[2,+∞)
{1,2,3,4,5}
1
Expert's answer
2021-09-30T23:39:56-0400

In the space of "X=E^1" let us find the complement of the following:

"(1,2],\ [2,+∞),\ \{1,2,3,4,5\}."


The complement of "(1,2]" is "(-\infty,1]\cup(2,+\infty)." The complement of of "[2,+\u221e)" is "(-\infty,2)." And the complement of "\{1,2,3,4,5\}" is "(-\infty,1)\cup(1,2)\cup(2,3)\cup(3,4)\cup(4,5)\cup(5,+\infty)."